Watson Company wants to raise capital for a planned expansion into a new market. The firm has 1 million shares of common equity with a par value (book value) of $1 and retained earnings of $30 million, its shares have a market value of $50 per share. It also has debt with a par or book value of $20 million, and 500,000 preferred shares outstanding.

You have collected the following information on Watson Company:

  • Watson has just paid a dividend of $3 and has expected dividend growth of 4.8% per year
  • Watson has a $20 million debt issue outstanding ($1000 par) with a 6% coupon rate. The debt has semi annual coupons and matures in five years. The bonds are selling at 95% of par
  • The company has a 40% tax rate
  • Watson also has 500,000 preferred shares outstanding. They are trading at $65 per share. They have a $75 par value and a dividend of $4 due in one year
  • The equity investors of Watson have a beta of 1.3. The T Bill rate is 5%, the market risk premium is 5%, and the return on the market is 10%.

A) What is your estimate of the average cost of equity capital?

B) What is the after tax cost of debt?

C) What is the cost of preferred capital?

D) What is the market value weight of the company’s common equity capital?

E) What is the market value weight of the company’s cost of debt capital?

F) What is the market value weight of the company’s cost of preferred equity capital?

G) What is the Weighted Average Cost of Capital (WACC) for Watson Company?
Answer #1
A] Cost of equity per 'constant dividend growth model' = Next expected dividend/Price +Growth rate = 3*1.048/50+0.048 = 11.09%
Cost of equity per CAPM = Risk free rate+beta*market risk premium = 5%+1.3*5% = 11.50%
Estimated average cost of equity = (11.09%+11.50%)/2 = 11.30%
B] Before tax cost of debt = YTM.
YTM, using an online financial calculator = 7.21%
After tax cost of debt = YTM*(1-t) = 7.21%*(1-40%) = 4.33%
C] Cost of preferred capital = 4/65 = 6.15%
D,E,F&G] Source of capital Market Value MV Weight Specific Cost WACC
Debt [20000000*95%] $      19,000,000 18.72% 4.33% 0.81%
Preferred stock [500000*65] $      32,500,000 32.02% 6.15% 1.97%
Common stock [1000000*50] $      50,000,000 49.26% 11.30% 5.56%
Total $    101,500,000 100.00% 8.34%
WACC = 8.34%